# Fuzzy Implicative Rules: A Unified Approach

### v1

@@ Abstract
Rule mining algorithms are one of the fundamental techniques in data mining for disclosing significant patterns in terms of linguistic rules expressed in natural language. In this paper, we revisit the concept of fuzzy implicative rule to provide a solid theoretical framework for any fuzzy rule mining algorithm interested in capturing patterns in terms of logical conditionals rather than the co-occurrence of antecedent and consequent. In particular, we study which properties should satisfy the fuzzy operators to ensure a coherent behavior of different quality measures. As a consequence of this study, we introduce a new property of fuzzy implication functions related to a monotone behavior of the generalized modus ponens for which we provide different valid solutions. Also, we prove that our modeling generalizes others if an adequate choice of the fuzzy implication function is made, so it can be seen as an unifying framework. Further, we provide an open-source implementation in Python for mining fuzzy implicative associative rules. We test the applicability and relevance of our framework for different real datasets and fuzzy operators.
Index Terms:
Knowledge discovery in databases (KDD), Rule-based models, Implicative rules, Association rules, Fuzzy logic, Fuzzy implication function.
@@ I Introduction
Data mining or Knowledge Discovery in Databases (KDD) is defined as the automatic extraction of patterns representing knowledge implicitly stored or captured in data [1]. Up to now, a wide variety of data mining techniques have been introduced and developed [2]. These techniques are usually divided into two types: exploratory and predictive. Exploratory data analysis focuses on searching relations between objects of a dataset (clustering, association rules, linguistic summaries…) whereas predictive data analysis aims to extract knowledge from discovered data with the intent to predict or classify unknown examples (classification, regression…). Within this context, rule-based algorithms have been one of the top choices for knowledge extraction because of their human-understandable output. Furthermore, with the emergence research area of eXplainable Artificial Intelligence (XAI) [3] they have become even more popular since they are highly interpretable models which are an appealing solution to provide an easily understandable representation of complex black-box models [4, 5].
In rule-based models, the output is given in terms of rules which are usually represented as $A\Rightarrow C$ in which $A$ is called the antecedent and $C$ is called the consequent. Although in these techniques the output has the same representation, depending on the knowledge to be captured we can find several rule mining techniques, just to mention some of them: (i) association rules: it is interested in capturing situations in which if $A$ is satisfied, then $C$ is likely to occur also [6]; (ii) classification rules: it aims to obtain a set of rules which effectively classify items into predefined classes [7]; (iii) sequential rules: it searches for rules that indicate that if some event(s) occurred, some other event(s) are likely to occur [8]; (iv) subgroup discovery: it aims to obtain rules with the most unusual statistical characteristics with respect to a property of interest [9]
[…]
@@
In this paper, we revisit the definition given in [11] for modeling fuzzy implicative rules and we contextualize it with respect to the advances in fuzzy operators made in recent years. Also, we generalize the four most used quality measures (coverage, support, confidence and weighted relative accuracy) to this new setting. Then, we study which properties should the pair of operators $(T,I)$ satisfy to obtain a coherent behavior of these measures. One of the key points of this study is to prove that for the support to be monotone with respect to the refinements of a rule we have to introduce an additional property of fuzzy implication functions related to the generalized modus ponens. Since up to our knowledge this property had not been previously introduced in the literature, we have studied it in the paper and we have proposed a list of families of operators that can be considered in this framework.
@@ II Preliminaries
$C(x_{1},y)x_{2}\geq C(x_{2},y)x_{1},\quad\text{for all }x_{1}\leq x_{2}\text{ and }y\in[0,1].$
In this case, $I_{C}$ is called a probabilistic implication.
@@ III-A Rule’s Modeling
$R^{S,L}_{j}:\text{IF }(X_{m_{1}}\text{ IS }LL_{m_{1}}^{n_{m_{1}}}\text{ AND }\dots\text{ AND }X_{m_{s}}\text{ IS }LL_{m_{s}}^{n_{m_{s}}})$
@@
Let $I$ be a fuzzy implication function, $T$ a t-norm and $R^{\tilde{S},\tilde{L}}_{\tilde{j}}$, $R^{S,L}_{j}\in\mathcal{R}^{I,T}$. We say that $R^{\tilde{S},\tilde{L}}_{\tilde{j}}$ is a refinement of $R^{S,L}_{j}$ if and only if $\tilde{j}=j$, $S\subsetneq\tilde{S}\text{ and }LL_{m}^{n_{m}}\in\tilde{L}$ for all $X_{m}\in S$. In this case, we denote it by $R^{S,L}_{j}\prec R^{\tilde{S},\tilde{L}}_{\tilde{j}}$.
@@
$T(\tilde{x},I(\tilde{x},y))\leq T(x,I(x,y)),\quad x,\tilde{x},y\in[0,1]\text{ with }\tilde{x}\leq x.$ (MTC)
@@
for $\text{for all }R^{S,L}_{j}\prec R^{\tilde{S},\tilde{L}}_{j},~{}R^{S,L}_{j},R^{\tilde{S},\tilde{L}}_{j}\in\mathcal{R}^{I,T}_{j}\text{ and }E^{d}\in E.$ Thus, $\sum_{d=1}^{n_{e}}\mu_{eval}^{R^{\tilde{S},\tilde{L}}_{j},E^{d}}\leq\sum_{d=1}^{n_{e}}\mu_{eval}^{R^{S,L}_{j},E^{d}}$.
@@ III-B Fuzzy implicative rules as a generalization of others
$\begin{array}[]{rcl}\mu_{Class_{j}}:D_{Y}&\longrightarrow&\{0,1\}\\ \tilde{j}&\longmapsto&\left\{\begin{array}[]{ll}1&\text{if }\tilde{j}=j,\\ 0&\text{if }\tilde{j}\not=j.\end{array}\right.\end{array}$
Then, it is obvious that when evaluating a rule for a certain example the only possible values for the consequent are 00 or $1$, so the only values of the fuzzy implication function that are being used are $I(x,1)$ and $I(x,0)$ for all $x\in[0,1]$. Since $I(x,1)=1$ for all $x\in[0,1]$, only the choice of the natural negation $N_{I}$ plays a role in this framework. However, if we take into account that in the end we are interested in the evaluation of the rule, if we choose a fuzzy implication function $I$ which satisfies (TC) with respect to $T$ we have that
@@ III-C Quality measures
Now, let us point out that the measure of fuzzy coverage is monotone with respect to the refinements, i.e., any refinement of a rule has a lower fuzzy coverage. This fact is also true for the fuzzy support if we impose that the pair ($T$,$I$) satisfies (MTC). This highlights again the importance of the newly introduced property. If we do not consider (MTC), the fuzzy support will not be necessarily monotone, and then it will not be adequate as a measure of generality.
@@ IV Solutions for adequate fuzzy operators
    $T(\tilde{x},I(\tilde{x},\tilde{y}))$ $=$ $T(\tilde{x},1)=\tilde{x}\leq y=T(y,1)$
@@
    $=$ $\left\{\begin{array}[]{ll}x&\text{if }x\leq y,\\ y&\text{if }x>y,\end{array}\right.$
@@
    $T(x,I_{T}(x,y))$ $=$ $\left\{\begin{array}[]{ll}T(x,1)&\text{if }x\leq y,\\ T(x,t^{-1}(t(y)-t(x)))&\text{if }x>y,\end{array}\right.$
    $=$ $\left\{\begin{array}[]{ll}x&\text{if }x\leq y,\\ t^{(-1)}(t(x)+t(y)-t(x))&\text{if }x>y,\end{array}\right.$
    $=$ $\left\{\begin{array}[]{ll}x&\text{if }x\leq y,\\ y&\text{if }x>y.\end{array}\right.$
@@
      $\geq$ $k^{-1}\left(\frac{k(\tilde{x})k(y)}{\tilde{x}}\right)\geq k^{-1}(k(\tilde{x}))$
@@
      $T_{k}(\tilde{x},I_{k}(\tilde{x},y))$ $=$ $k^{-1}\left(\frac{k(\tilde{x})k(y)}{\tilde{x}}\right)$
@@
  • If $I_{C}$ is a probabilistic implication, then it satisfies (MTC) with respect to the product t-norm. Thus, $(T_{\bm{P}},I_{C})$ is an adequate pair for fuzzy implicative rule mining.
  • If $k$ is a $k$-generator with $k(\tilde{x})\cdot x\leq k(x)\cdot\tilde{x}$ for all $0\leq\tilde{x}\leq x\leq 1$ then $(T_{k},I_{k})$ where $I_{k}$ is the $k$-generated implication and $T_{k}$ the t-norm generated by $k$ as its multiplicative generator satisfy (MTC). In this case, if we consider $x=1$ then we have $k(\tilde{x})\leq k(1)\cdot\tilde{x}=\tilde{x}=\text{id}(\tilde{x})$ for all $0\leq\tilde{x}\leq 1$ and the pair $(T_{k},I_{k})$ also satisfies (TC). In particular, $k(0)=0$ in this case. Then, $(T_{k},I_{k})$ is an adequate pair for fuzzy implicative rule mining under these conditions. For instance, let us consider the following $k$-generators:
@@
  • The point (v)-Proposition 26 discloses that the family of $T$-power invariant implications does not provide interesting solutions when studying the property (MTC). Indeed, any solution of this property returns zero when applying the generalized modus ponens. Thus, for the moment there are no solutions if we want to use fuzzy implication functions which are invariant with respect to hedges modeled through powers of t-norms.
  • Finally, if we consider any t-norm and the fuzzy implication function in Eq. (1) then we obtain a pair of operators which is adequate for fuzzy implicative rule mining and the modeling is equivalent to fuzzy conjunctive rules.
Summarizing the above discussion, in the following proposition we point out five different pairs of fuzzy implication function and t-norm that are adequate for fuzzy implicative rule mining. More specifically, in Table I we propose four concrete examples of adequate pairs which are the ones considered in our experimental results.
@@
Once we have several options for the selection of the pair of operators $(T,I)$ it is interesting to look at the structure of the evaluation of the antecedent and the rule for different pairs. In Proposition 26 we can see the expression of the truth value of the evaluation of the rule when the generalized modus ponens is used for each case. Notice that even though fuzzy implication functions are non-commutative, it may happen that the expression given by the generalized modus ponens is commutative. Indeed, this is the case for $R$-implications and also probabilistic implications that use a commutative copula. The user might like to avoid these options if the purpose is to obtain directional rules, in which it is significantly different the permutation of antecedent and consequent. That is why in the last two rows in Table I we have considered pairs of operators that result in non-commutative evaluations of the corresponding rules by considering a non-commutative copula for a probabilistic implication and a $k$-implication.
@@
Product t-norm
$T_{\bm{P}}(x,y)=xy$.
Probabilistic fuzzy implication based on
$(C_{\lambda}^{H})_{1,0}$ copula
$I_{\bm{(C_{\lambda}^{H})_{1,0}}}(x,y)=$
$\par\left\{\begin{array}[]{ll}1&\text{ if }x=0,\\ \frac{y}{x}-\frac{(1-x)y}{x(\lambda+(1-\lambda)(1-x+xy))}&\text{ if }x>0,\end{array}\right.$
with $\lambda\in[0,2]$. Considered value $\lambda=0.01$.
@@
with $\lambda\in(-\infty,0)$. Considered value $\lambda=-10$.
V Case Studio: Fuzzy implicative association rules
In this section, we apply fuzzy implicative rules to the specific problem of mining association rules. To achieve this, we have implemented an algorithm similar to Apriori [27], incorporating redundancy pruning as described in [28]. This algorithm identifies frequent itemsets based on a minimum fuzzy coverage threshold and then mines association rules by considering minimum thresholds for fuzzy support and confidence. Unlike other approaches, in our framework, support is generally not symmetric. This asymmetry has been taken into account when deriving the final rules. The Python implementation can be found in the following open repository https://github.com/rferper/FIRM, in which we allow the custom selection of fuzzy partitions and operators. Since in Section III-B we have proved that our framework generalizes the crisp and other fuzzy perspectives, instead of comparing with other implementations (see for instance [23, 29]), we will only compare different choices of fuzzy operators. This makes the comparison more robust, since we ensure that the fuzzification process is exactly the same in all evaluations.

### v2

@@ Abstract
Rule mining algorithms are one of the fundamental techniques in data mining for disclosing significant patterns in terms of linguistic rules expressed in natural language. In this paper, we revisit the concept of fuzzy implicative rule to provide a solid theoretical framework for any fuzzy rule mining algorithm interested in capturing patterns in terms of logical conditionals rather than the co-occurrence of antecedent and consequent. In particular, we study which properties should satisfy the fuzzy operators to ensure a coherent behavior of different quality measures. As a consequence of this study, we introduce a new property of fuzzy implication functions related to a monotone behavior of the generalized modus ponens for which we provide different valid solutions. Also, we prove that our modeling generalizes others if an adequate choice of the fuzzy implication function is made, so it can be seen as an unifying framework. We test the applicability and relevance of our framework for different real datasets and fuzzy operators.
@@ I Introduction
Data mining or Knowledge Discovery in Databases (KDD) is defined as the automatic extraction of patterns representing knowledge implicitly stored or captured in data [1]. Up to now, a wide variety of data mining techniques have been introduced and developed [2]. These techniques are usually divided into two types: exploratory and predictive. Exploratory data analysis focuses on searching relations between objects of a dataset (clustering, association rules, linguistic summaries…), whereas predictive data analysis aims to extract knowledge from discovered data with the intent to predict or classify unknown examples (classification, regression…). Within this context, rule-based algorithms have been one of the top choices for knowledge extraction because of their human-understandable output. Furthermore, with the emergence of the research area of eXplainable Artificial Intelligence (XAI) [3], they have become even more popular since they are highly interpretable models, which are an appealing solution to provide an easily understandable representation of complex black-box models [4, 5].
In rule-based models, the output is given in terms of rules which are usually represented as $A\Rightarrow C$ in which $A$ is called the antecedent and $C$ is called the consequent. Although in these techniques the output has the same representation, depending on the knowledge to be captured we can find several rule mining techniques, just to mention some of them: (i) association rules: it is interested in capturing situations in which if $A$ is satisfied, then $C$ is likely to occur also [6]; (ii) classification rules: it aims to obtain a set of rules which effectively classify items into predefined classes [7]; (iii) sequential rules: it searches for rules that indicate that if some event(s) occurred, some other event(s) are likely to occur [8]; (iv) subgroup discovery: it aims to obtain rules with the most unusual statistical characteristics with respect to a property of interest [9], etc.
Despite the representation of the rules as logical conditionals, they are usually interpreted and evaluated as the co-occurrence of antecedent and consequent. Although this assumption is quite conventional in the literature, it has also received some criticism [10, 11]. Indeed, the discrepancy between the linguistic representation of the rules and their underlying mathematical modeling can be confusing for the expert. For instance, if one considers the rule $\text{Smoking}\Rightarrow\text{Respiratory Problems}$ it might seem unnatural that the support of this rule is not affected by its direction, because there exists a lot of respiratory problems that are not associated with smoking. An alternative to overcome this drawback is to consider a more logic-oriented approach in which the conditional in the rule is interpreted as a logical implication, in this case normally the term “implicative rules” is used. In the crisp setting, the authors in [10] proposed to interpret the logical conditional through a measure called “conviction” as an alternative to confidence which is based on conditional probability. However, in the fuzzy setting, fuzzy implication functions can be directly used. Indeed, in [11] the author introduced implication-based fuzzy association rules by using a t-norm for modeling conjunction, a fuzzy implication function to model the logical conditional, and the generalized modus ponens as inference schema. Nonetheless, no further development of this framework or a thorough theoretical study of the proposal was made.
@@
In this paper, we revisit the definition given in [11] for modeling fuzzy implicative rules and we contextualize it with respect to the advances in fuzzy operators made in recent years. We generalize the four most used quality measures (coverage, support, confidence, and weighted relative accuracy) to this new setting. Then, we study which properties the pair of operators $(T,I)$ should satisfy to obtain a coherent behavior of these measures. One of the key points of this study is to prove that for the support to be monotone with respect to the refinements of a rule, we have to introduce an additional property of fuzzy implication functions related to the generalized modus ponens. To our knowledge, this property had not been previously introduced in the literature, so we have studied it in the paper, and we have found several solutions. Also, we prove that not every choice of pair of operators $(T,I)$ models directional rules; in fact, we show that the most common fuzzy implication functions are equivalent to the use of t-norms in this setting, so we propose different pairs of operators that yield information beyond what t-norms can provide.
@@ II Preliminaries
$C(x_{1},y)x_{2}\geq C(x_{2},y)x_{1},$ (1)
for all $x_{1}\leq x_{2}\text{ and }y\in[0,1]$. In this case, $I_{C}$ is called a probabilistic implication.
@@ III-A Rule’s Modeling
Let $I$ be a fuzzy implication function, $T$ a t-norm and $R^{\tilde{S},\tilde{L}}_{\tilde{j}}$, $R^{S,L}_{j}\in\mathcal{R}^{I,T}$. We say that $R^{\tilde{S},\tilde{L}}_{\tilde{j}}$ is a refinement of $R^{S,L}_{j}$ if and only if $\tilde{j}=j$, SS~ $\subsetneq\tilde{S}\text{ and }LL_{m}^{n_{m}}\in\tilde{L}$ for all $X_{m}\in S$. In this case, we denote it by $R^{S,L}_{j}\prec R^{\tilde{S},\tilde{L}}_{\tilde{j}}$.
@@
[…]
@@
$\text{for all }R^{S,L}_{j}\prec R^{\tilde{S},\tilde{L}}_{j},~R^{S,L}_{j},R^{\tilde{S},\tilde{L}}_{j}\in\mathcal{R}^{I,T}_{j}\text{ and }E^{d}\in E.$ Thus, $\sum_{d=1}^{n_{e}}\mu_{eval}^{R^{\tilde{S},\tilde{L}}_{j},E^{d}}\leq\sum_{d=1}^{n_{e}}\mu_{eval}^{R^{S,L}_{j},E^{d}}$.
@@ III-B Fuzzy implicative rules as a generalization of others
[…]
Then, it is obvious that when evaluating a rule for a certain example the only possible values for the consequent are 0 or $1$, so the only values of the fuzzy implication function that are being used are $I(x,1)$ and $I(x,0)$ for all $x\in[0,1]$. Since $I(x,1)=1$ for all $x\in[0,1]$, only the choice of the natural negation $N_{I}$ plays a role in this framework. However, if we take into account that in the end we are interested in the evaluation of the rule, if we choose a fuzzy implication function $I$ which satisfies (TC) with respect to $T$ we have that
@@ III-C Quality measures
Now, let us point out that the measure of fuzzy coverage is monotone with respect to the refinements, i.e., any refinement of a rule has a lower fuzzy coverage. This fact is also true for the fuzzy support if we impose that the pair ($T$,II) satisfies (MTC). This highlights again the importance of the newly introduced property. If we do not consider (MTC), the fuzzy support will not necessarily be monotone, and then it will not be adequate as a measure of generality. Indeed, it would be inadmissible to allow that rules with fewer conditions in the antecedent have less support than more complex rules derived from them, as it contradicts human common sense. Also, without this property, support pruning would not be possible, which clearly affects any optimization process.
@@ IV Solutions for adequate fuzzy operators
    $)\displaystyle T(\tilde{x},I(\tilde{x},\tilde{y}$)) $=$ $T(\tilde{x},1)=\tilde{x}\leq y=T(y,1)$
@@
    $T_{k}(x,I_{k}(x,y))=k^{-1}\left(k(x)\cdot\min\left\{\frac{k(y)}{x},1\right\}\right).$ (7)
@@
[…]
@@
    $T(x,I_{T}(x,y))$ $=$ $\left\{\begin{array}[]{ll}\hskip-2.84526ptT(x,1)&\hskip-2.84526pt\text{if }x\leq y,\\ \hskip-2.84526ptT(x,t^{-1}(t(y)-t(x)))&\hskip-2.84526pt\text{if }x>y,\end{array}\right$.
    $=$ $\left\{\begin{array}[]{ll}\hskip-2.84526ptx&\hskip-5.69054pt\text{if }x\leq y,\\ \hskip-2.84526ptt^{(-1)}(t(x)+t(y)-t(x))&\hskip-5.69054pt\text{if }x>y,\end{array}\right$.
[…]
@@
      $\geq$ $)\displaystyle k^{-1}\left(\frac{k(\tilde{x})k(y)}{\tilde{x}}\right)\geq k^{-1}(k(\tilde{x}$))
@@
[…]
@@
  • If $I_{C}$ is a probabilistic implication, then it satisfies (MTC) with respect to $T_{\bm{P}}$. Thus, $(T_{\bm{P}},I_{C})$ is an adequate pair for fuzzy implicative rule mining.
[…]
@@
  • The point (v)-Proposition 26 discloses that the family of $T$-power invariant implications does not provide interesting solutions when studying the property (MTC). Indeed, any solution of this property returns zero when applying the generalized modus ponens. Thus, for the moment, there are no solutions if we want to use fuzzy implication functions that are invariant with respect to hedges modeled through powers of t-norms.
  • Finally, if we consider any t-norm and the fuzzy implication function in Eq. (2), then we obtain a pair of operators that is adequate for fuzzy implicative rule mining, and the modeling is equivalent to fuzzy conjunctive rules.
Once we have several options for the selection of the pair of operators $(T,I)$ it is interesting to look at the structure of the evaluation of the antecedent and the rule for different pairs. In Proposition 26 we can see the expression of the truth value of the evaluation of the rule when the generalized modus ponens is used for each case. Notice that even though fuzzy implication functions are non-commutative, it may happen that the expression given by the generalized modus ponens is commutative. Indeed, this is the case for $R$-implications and also probabilistic implications that use a commutative copula. What is more, any choice of $R$-implication, one of the most popular fuzzy implication functions in fuzzy inference, is equivalent to using the minimum t-norm, and then it does not provide useful and different information for our framework. Thus, the user might like to avoid these options if the purpose is to obtain directional rules, in which the permutation of antecedent and consequent is significantly different. Then, we have seen that it may be hard to find a pair of operators $(T,I)$ so the value $T(x,I(x,y))$ fulfills all the desired properties. Indeed, although probabilistic implications are an option, it is not so easy to find a non-commutative copula fulfilling Eq. (1). Besides, $k$-implications reduce the value of that expression to 1 in a certain subdomain of $0,1]^{2}$ (see Eq. (7)). Nonetheless, it is not hard to propose a pair $(T,I)$ that fulfills all the conditions and, in contrast with those operators already introduced in the literature, it may be more tailored for a specific problem. For instance, we propose the following fuzzy implication function.
Proposition 27.
Let $p\in(0,+\infty)$ and $I_{p}:[0,1]^{2}\to[0,1]$ defined as follows:
$I_{p}(x,y)=1-x+xy^{p},\quad\text{for all }x,y\in[0,1].$ (13)
Then $I_{p}$ is a fuzzy implication function with $T_{\bm{LK}}(x,I_{p}(x,y))=xy^{p}$ and the pair $(T_{\bm{LK}},I_{p})$ fulfills (TC) and (MTC).
Proof.
First, we prove that $I_{p}$ is a fuzzy implication function. Indeed, $I_{p}(0,0)=I_{p}(1,1)=0$, $I_{p}(1,1)=1$ and
$x_{1}\leq x_{2}\Rightarrow 1+x_{1}(y^{p}-1)\geq 1+x_{2}(y^{p}-1),$
$y_{1}\leq y_{2}\Rightarrow 1+x(y_{1}^{p}-1)\leq 1+x(y_{2}^{p}-1),$
so we obtain $_{p}I(x_{1},y)\geq I_{p}(x_{2},y)$ and $I_{p}(x,y_{1})\leq I_{p}(x,y_{2})$.
$T_{\bm{LK}}(x,I_{p}(x,y))=\max\{x+1-x+xy^{p}-1,0\}=xy^{p}.$
Now, since $xy^{p}\leq x$ for all $x,y\in[0,1]$ and $x_{1}y^{p}\leq x_{2}y^{p}$ we know that the pair $(T_{\bm{LK}},I_{p})$ fulfills (TC) and (MTC). ∎
With the pair of operators in Proposition 27, we can adjust the importance of the consequent by modifying the parameter $p$; if $p>1$, we shrink its contribution, whereas when $p<1$, we inflate it. Moreover, the functional form of the operators is simple and easily implementable. We see in Section V-A how this non-commutative behavior is a key point when trying to detect directional patterns.
Summarizing the above discussion, in the following proposition, we point out six different pairs of fuzzy implication functions and t-norms that are adequate for fuzzy implicative rule mining. More specifically, in Table I we propose four concrete examples of adequate pairs, which are the ones considered in our experimental results.
@@
  • $(T_{\bm{LK}},I_{p}))$, where $I_{p}$ is the fuzzy implication function in Eq. (13).
@@
with $\lambda\in(-\infty,0)$.
Łukasiewicz t-norm
$T_{\bm{LK}}(x,y)=\max\{x+y-1,0\}$.
$I_{p}(x,y)=1-x+xy^{p}.$
V Experiments
In this section, we discuss the role of fuzzy implicative rules in the specific problem of mining association rules, which we use as an illustrative case to demonstrate the utility of the framework. For simplicity, we rely on the well-known exhaustive Apriori algorithm [27], enhanced with redundancy pruning as described in [28]. Apriori identifies frequent itemsets based on a minimum fuzzy coverage threshold and subsequently derives association rules by applying minimum thresholds for fuzzy support and confidence. Unlike classical approaches, in our framework support is not necessarily symmetric, and this asymmetry is explicitly taken into account when formulating the final rules. Consequently, for each frequent itemset, we must consider all possible directional antecedent–consequent splits. Accordingly, the rule-generation cost increases with the frequent itemset size, yet empirically, it did not dominate the runtime for the datasets evaluated. The frequent itemset generation cost increases exponentially with the number of items, just as classical Apriori. Nonetheless, coverage and support thresholds significantly lower the computation time. The Python implementation is available in the open repository https://github.com/rferper/FIRM , where we allow the user to select custom fuzzy partitions and operators. Moreover, as we proved in Section III-B, our framework generalizes other fuzzy approaches, so we use our implementation to compare with existing fuzzy association mining tools such as the lfl package [29]. For comparison with crisp association rules, we use the arules package [30]. In this case, discretization is performed using the same cuts as the fuzzy sets to ensure a fair evaluation.
V-A Mining directional patterns
This section presents a controlled experiment illustrating how fuzzy association rules may or may not capture directional dependencies. We first generated a synthetic bivariate dataset designed to encode an asymmetric relation. Variable $A$ was sampled uniformly on $[0,100]$. Conditional on $A$, variable $B$ followed
$B\sim\begin{cases}\mathcal{U}(0,100),&A\leq 70,\\[4.0pt] \mathcal{U}(70,100),&A>70,\end{cases}$
so that large values of $A$ enforce correspondingly large values of $B$. This mechanism creates a structured gap in the joint distribution. In total, we generated 1000 points and the resulting scatter plot is shown in Figure 1.
Figure 1: Synthetic dataset illustrating an induced asymmetric dependency between two variables $A$ and $B$.
From this construction, it follows that if $A$ is High then $B$ is High, but not necessarily the other way around. Moreover, no other rules are expected to hold significance.
We now examine how support and confidence evolve for operator pairs exhibiting non-commutative behavior. Figure 2 shows that, for $(T_{\lambda}^{\bm{SS}},I^{\bm{SS}}_{k_{\lambda}})$, support and confidence remain nearly invariant, with both rules showing very similar support. Although $A$ High $\to$ $B$ High achieves slightly higher confidence than the reverse, the difference is negligible. Hence, these operators, despite being non-commutative, are too restrictive to capture the intended directional pattern. In contrast, for $(T_{\bm{LK}},I_{p})$ the behavior changes substantially. For $p<1$, support and confidence diverge clearly between the two rules, correctly highlighting the true dependency. For $p>1$, however, the opposite occurs: the rule $B$ High $\to$ AA High gains inflated confidence due to overemphasizing the antecedent, which is frequently satisfied when $B$ is High.
(a) $(T_{\lambda}^{\bm{SS}},I^{\bm{SS}}_{k_{\lambda}})$.
(b) $(T_{\bm{LK}},I_{p})$.
Figure 2: Evolution of support and confidence for the rules AA High $\to$ $B$ High and $B$ High $\to$ AA High as a function of the parameters $\lambda$ and $p$ of the pairs $(T_{\lambda}^{\bm{SS}},I^{\bm{SS}}_{k_{\lambda}})$ and $(T_{\bm{LK}},I_{p})$, respectively.
For comparison, we also computed support and confidence for fuzzy association rules using both the minimum and product t-norms, as well as for the crisp case. The results are reported in Table II. Across all three settings, the values are nearly identical and do not distinguish between the two rules. As expected, support remains identical for both rules in all cases.
TABLE II: Support and confidence values for the rules AA High $\to$ $B$ High and $B$ High $\to$ AA High under different operator choices: product t-norm with Yager implication $(T_{P},I_{Y})$, Łukasiewicz t-norm with Łukasiewicz implication $(T_{LK},I_{LK})$, and the crisp case.
AA High $\to$ $B$ High $B$ High $\to$ AA High
$T_{\bm{P}}$ (0.23,0.63) (0.23,0.62)
$T_{\bm{M}}$ (0.24,0.64) (0.24,0.63)
Crisp (0.23,0.64) (0.23,0.62)
In summary, we have shown that for recovering the true asymmetric dependency, we require operators with non-symmetric support and explicit control over the consequents’ contribution. Both properties become available through the introduction of fuzzy implication functions.
V-B Similarity between models